make x subject of the formula ux^2+t=5t-2x^2
Answer by Theo(13342)

start with ux^2+t = 5t-2x^2

subtract t from both sides of the equation to get ux^2 = 5t-2x^2-t

combine like terms to get ux^2 = 4t-2x^2

add 2x^2 to both sides of the equation to get ux^2+2x^2 = 4t

factor our x^2 from ux^2+2x^2 to get x^2*(u+2) = 4t

divide both sides of the equation by (u+2) to get x^2 = 4t/(u+2)

take the square root of both sides of the equation to get x = +/- sqrt(4t/(u+2))

that should be your solution.

you have solved for x in terms of the other variables in the equation.